250 cm3 of 0.2M of k2so4 mixed with 250 cm3 of 0.2M of kcl.Find the molar concentration of k+ ions in solution

Grade:12th pass

2 Answers

Aditya Gupta
4 years ago
hello student, kindly follow the link below:
feel free to ask any doubt if u have any :))
Vikas TU
4 years ago
first calculate the volume . 
if u will calculate the volume u will get it as 500 cc .. 
after that find the total no.of K+ ions.
then calculate by the following formula 
volume*molarity*n factor 
and acc to the question by calculating 
u will get it as 
K+ = 250*0.2*1 + 250*0.2*2 = 150
the molarity will be 0.2 in each case here .. 
then for finding the final answer apply the following formula 
conc of K+ = no.of K+/volume
by applying the above formula u will get 
 150/500 = 0.3M
which will be the answer 
HOPE U WILL GET HOW TO SOLVE SUCH QUESTIONS 
ALL THE BEST